OpenStudy (anonymous):

integration of 1/x^4(x^2+1)^1/2

OpenStudy (anonymous):

\[\int\limits \frac{ dx }{ x ^{4}\sqrt{x ^{2}+1} }\]

hartnn (hartnn):

did you try to put x = tan u ?

OpenStudy (anonymous):

you mean \[\tan \theta \] if so yes i tried

hartnn (hartnn):

yeah, theta is not so convenient to write, so i used 'u' , or we can use 't' x = tan t will simplify things, what is your new intergral in terms of t ?

OpenStudy (anonymous):

\[\int\limits\frac{ \sec ^{2}\theta }{ \tan^4\theta \sec \theta }d \theta\]

OpenStudy (anonymous):

it was \[\int\limits\frac{ \sec ^{2}\theta }{ \tan^4\theta \sqrt{\tan^2 \theta+1} }d \theta\]

OpenStudy (anonymous):

i used \[\sqrt{x^2+1} =\sec \theta \]

hartnn (hartnn):

yeah so, sec t cot^4 t since we have cot t lets try to bring csc t sec t * cot t = 1/cos t * cos t / sin t = csc t so, it'll be cot^3 t csc t dt

hartnn (hartnn):

see if you get that ^

OpenStudy (anonymous):

The attachment shows how Mathematica does it.

hartnn (hartnn):

we don't need full solutions i think, if mostafa wants to learn...

OpenStudy (anonymous):

yes and no

OpenStudy (anonymous):

i will try to get what you said hartnn

OpenStudy (anonymous):

finally i got it thanks

hartnn (hartnn):

good! :)